OpenStudy (anonymous):

Cat Years 1 3 4 6 7 Human Years 13 33 43 63 73 What is the meaning of the slope for this data? A. 20 years of a cat's life is equivalent to one year of a human's life. B. One year of a cat's life is equivalent to about 10 years of a human's life. C. One year of a cat's life is equivalent to 5 years of a human's life.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Notice that for human years, 33-13=20 For cat years: 3-1=2 Thus the ratio is 20/2=10/1 meaning that for 10 years of human life, at cat is only 1 year old

OpenStudy (anonymous):

oh okay i get now. thanks